# Randomized block Krylov method for approximation of truncated tensor SVD

### v2

@@ Abstract
[inst1]organization=Lab of Machine Learning and Knowledge Representation, Innopolis University, 420500,city=Innopolis, country=Russia, s.ahmadiasl@innopolis.ru
@@ 1 Introduction
Similar to SVD, which is used to analyze matrices, tensor decompositions are mathematical tools that generalize matrix decompositions (e.g., SVD, PCA) to higher-order tensors. They aim to represent a tensor as a combination of simpler, lower-dimensional components. The most common tensor decompositions include: Canonical polyadic decomposition [15], Tucker decomposition [16], tensor train decomposition [17], tensor ring decomposition [18] and tensor SVD (T-SVD) [19]. Tensor decompositions have numerous applications across various domains, such as signal processing to machine learning, tensors are indispensable tools in modern data science. The randomized block Krylov subspace method was proposed in [20] and was shown to be a very efficient technique for the low-rank approximation of matrices. Motivated by these results, it was extended to several tensor decompositions. For example, it was previously utilized in [21] and [22] for Tucker and TT decompositions, respectively. It is known that the randomized block Krylov subspace method can better capture the range of a matrix compared to the classical randomized algorithms, and this leads to obtaining better low-rank approximations. It is worth highlighting that randomized algorithms are applicable for computing low-rank approximations of data tensors that have low-rank structures, such as images and videos.
In this paper, we propose to use the randomized block Krylov method for the computation of the truncated Tensor SVD (T-SVD). It is another extension of the method proposed in [20] to tensors. To the best of our knowledge, this is the first paper discussing the possibility of using the randomized block Krylov method for the T-SVD model. A detailed theoretical analysis of the proposed algorithm is discussed, and they are supported by simulation results. An application of the proposed randomized method to image completion is given. Our simulation results show the effectiveness of the proposed randomized algorithm for the synthetic and real-work datasets.
@@ 2 Preliminary concepts
This section presents some definitions and lemmas that will be used in the sequel. We denote tensors with calligraphic letters and matrices with capital letters. The set of real and natural numbers denote by $\mathbb{R}$ and $\mathbb{N}$, respectively. Let $O$, and $\mathcal{O}$ stand for a zero matrix, and a zero tensor, respectively.
@@
where $\mathcal{U}\in\mathbb{R}^{n_{1}\times n_{1}\times n_{3}}$ is an orthogonal tensor, $\mathcal{S}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ is an f-diagonal tensor (i.e., each frontal slice is a diagonal matrix), $\mathcal{V}\in\mathbb{R}^{n_{2}\times n_{2}\times n_{3}}$ is an orthogonal tensor. The number of nonzero tubes in the middle tensor $\mathcal{S}$ is called tubal rank of $\mathcal{X}$ and is denoted by TrankT-rank. The truncated T-SVD of tubal rank $R$ for the tensor $\mathcal{X}$} is
@@ 1 $\overline{{\mathcal{X}}}=\mathtt{fft}\left({{\mathcal{X}},[],3}\right)$;
1 $\overline{{\mathcal{X}}}=\mathtt{fft}\left(\overline{X},[],3\right)$;
@@ 3 Proposed randomized block Krylov method for approximate truncated T-SVD
The randomized block Krylov subspace algorithm is an efficient method for approximating large-scale matrix computations, such as eigenvalue problems or low-rank matrix approximations. It combines randomization techniques with the power of Krylov subspace methods to achieve high accuracy with reduced computational cost. Although a classical randomized algorithm applies power iterations $q\geq 1$ to enhance the decay of singular values, it does not use the results from each step of the iteration. Since it only uses the final result, $(\mathcal{X}*\mathcal{X}^{\top})^{q}*\mathcal{X}*\mathcal{B}$, the method generally cannot closely approximate the largest k singular values. This captures a much richer spectral range, leading to faster, super-linear convergence and higher accuracy for the same number of tensor operations. Let us discuss this precisely. The classical randomized algorithm with the power iteration technique for truncated T-SVD is presented in Algorithm 4. As can be seen, it uses the last iteration $(\mathcal{X}*\mathcal{X}^{\top})^{q}*\mathcal{X}*\mathcal{B}$ of the power iteration loop for computing an orthonormal basis of the tensor range in Line 7. The proposed randomized algorithm looks for a better matrix to capture the range of the matrix and do this, it uses all iterations $(\mathcal{X}*\mathcal{X}^{\top})^{i}*\mathcal{X}*\mathcal{B},\,\,i=1,2,\ldots,q$, which leads to better capturing the range of the tensor and better approximation results. This process is summarized in Algorithm 5. So, the main difference between the proposed randomized algorithm and the one proposed in [30] is how they compute an orthonormal basis for the range of a tensor to be used for the projection stage within the randomization framework. We extensively examine this idea with detailed theoretical analyses.
@@ 11 $\widehat{\mathcal{U}}=\mathcal{Q}*\mathcal{U}_{c}$;
$\widehat{\mathcal{U}}_{R}=\widehat{\mathcal{U}}(:,1:R,:),\widehat{\mathcal{V}}_{R}=\mathcal{V}_{c}(:,1:R,:),\widehat{\mathcal{S}}_{R}=\mathcal{S}_{c}(1:R,1:R,:)$
@@ 4 Theoretical results
In this section, we present a detailed theoretical analysis of the proposed randomized subspace block Krylov method. A key point needs to be highlighted is that the proposed algorithm indeed applies the randomized block Krylov method on each frontal slice of the tensor in the Fourier domain. This facilitates deriving the upper bound approximation obtained by the proposed algorithm.
@@
Since spectral norm on matrices is unitarily invariant and $\mathtt{bcirc}$ of an orthogonal tensor is an orthogonal matrix, we have
@@
Let $\mathcal{A}\in\mathbb{R}^{n\times n\times n_{3}}$ be a T-symmetric tensor and $\mathcal{C}$ be a real tensor with adequate dimension. If $\mathcal{A}$ is a T-positive semidefinite tensor, then $\mathcal{C}^{\top}*\mathcal{A}*\mathcal{C}$ is a T-positive semidefinite tensor.
@@
Since each term of $\mathcal{D}$ is T-positive semidefinite, it can be easily show that $\mathcal{D}$ is also semidefinite. So, from Lemma 11, we have
@@
where $\tilde{\mathcal{U}}$ and $\tilde{\mathcal{V}}$ are orthogonal tensors and $\tilde{\mathcal{S}}$ is an f-diagonal tensor, in which the entries on the first frontal slice are the singular values of the tensor $\mathcal{S},(\mathcal{S}*\mathcal{S}^{\top})*\mathcal{S},\ldots,(\mathcal{S}*\mathcal{S}^{\top})^{q}*\mathcal{S}]$. In other words, they are square root of eigenvalues of the following tensor
@@ 5 Experimental results
The proposed randomized algorithm and the baseline (Algorithm 4) are applied to the mentioned data tensor with the parameters P=5,q=2P=5,\,q=2 and $R=45$. The relative errors and running times of the algorithms are reported in Figures 1-3. We should point out that in Algorithm 5, in Line 7, the factor $\mathcal{Q}$ was truncated to $k+p$ term as our simulations showed that there is no difference between truncation and the full factor for this synthetic data, and our reported running times are for this case. If we consider the full factor $\mathcal{Q}$, the running time of Algorithm 5 will be higher than Algorithm 4. What we see from our numerical experiments that the proposed randomized algorithm can provide estimates with better accuracy compared to the baseline with a little higher running time. As was discussed in Section 3, the superior performance of Algorithm 5, is due to the fact that it can capture a much richer spectral range, leading to faster, superlinear convergence and higher accuracy for the same number of tensor operations.
We also made a sensitivity analysis on the oversampling and power iteration parameters. Our observation was that for tensors with frontal slices with fast singular value decay, the small oversampling parameter, e.g. $p=5,10,20$ is sufficient, often independent of the target rank $R$. Also, for the power iteration parameter, the small integers, e.g., $q=0,1,2,3$ were sufficient and rarely exceeded $q=4$ due to numerical stability concerns.
@@
(Image compression) In this example, we consider the application of the proposed randomized block Krylov subspace method for compressing color images. We consider three color images “peppers”, “baboon” and “house” depicted in Figure 4. All images have size $256\times 256\times 3$. Then, the proposed randomized approach and the benchmark Algorithm 4 were applied to the three images using P=5,q=2P=5,\,q=2 and the tubal rank $R=25$. The reconstructed images computed by them are shown in Figure 4. The proposed algorithm can achieve marginally better results than the baseline, while its running time is a little higher due to dealing with a block of matrices (the tensor $\mathcal{K}$ in Algorithm 5) instead of only one single matrix (the tensor $\mathcal{B}$ in Line 7 of Algorithm 4). This requires extra computations. The simulation results show that the proposed randomized algorithm could provide better results than the baseline. The sensitivity analysis on the oversampling and power iteration parameters were the same as the first example. More precisely, for the oversampling parameter, we achieved smooth improvement for 1 to 10. For the power iteration from $q=1$ to, $q=2,$ we saw a significant improvement, while for the $q\geq 2$ the improvement was not so much.
@@
(Image completion) In this experiment, we apply the proposed randomized block Krylov subspace method for the data completion task. We use the methodology proposed in [10, 37], where a low-rank approximation is required in the completion process. For the paper to be self-contained, let us briefly describe it. The image completion can be formulated as an optimization problem:
@@
${{\bf P}_{\mathcal{T}}}\left({\mathcal{X}}\right)=\left\{\begin{array}[]{l}{x_{{i_{1}},{i_{2}},\ldots,{i_{N}}}}\,\,\,\,\,\,\left({{i_{1}},{i_{2}},\ldots,{i_{N}}}\right)\in\mathcal{T}.\\ 0\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm Otherwise}.\end{array}\right$.
@@
${{\bf P}_{{\mathcal{T}}}}({{\mathcal{C}}})={{\bf P}_{{\mathcal{T}}}}({{\mathcal{M}}})$ (27)
@@
Three completion methods as baselines, including: Bayesian tensor CPD (Canonical Polyadic decomposition) completion [^3] https://github.com/qbzhao/BCPF [38], tensor ring completion [^4] https://github.com/HuyanHuang/Tensor-completion-via-tensor-ring-decomposition [39] and tensor train completion [^5] https://github.com/zhaoxile/ [40]. The comparison was made for the “kodim03” image and for high missing ratios. The recovered images by these algorithms are shown in Figure 9. For all of these algorithms, the default parameters and tensor ranks are used in the demo files of their GitHub pages. Our experiments show the method is highly effective for image completion, performing robustly under both random and structured missing patterns. Comparing our results with the baselines for 95%, and 98% shows significant improvement. The Bayesian tensor CPD and tensor train completion were very slow, but the tensor ring completion algorithm was faster. The running time of our algorithm was comparable to the tensor ring completion, but it was approximately 4-5 times faster.
@@
Figure 9: The completion results obtained by a) Bayesian tensor CPD (Canonical Polyadic decomposition) completion [38], b) tensor ring completion [39] and c) tensor train completion [40] for images with high missing ratios, for example 3.
@@
For this experiment, we use the Kodak dataset [^6] https://r0k.us/graphics/kodak/, comprising 23 images of size either $768\times 512\times 3$ or $512\times 768\times 3$. For consistency, all images were resized to $256\times 256\times 3$. We remove 80% of pixels in all images and using tubal tensor rank $R=40$, the oversampling parameter, $P=10$ and the power iteration parameter $q=2$. We use 100 iterations of the completion algorithm for the image recovery process. The PSNRs of the recovered images for the whole Kodak dataset are shown in Figure 10. This experiment showed the feasibility of the completion algorithm empowered by the proposed randomized block Krylov method. On average, completion of each image took 13 seconds; however, the algorithm’s computation can be significantly accelerated on GPU devices. Furthermore, when processing a batch of images, the completion process can be executed in parallel for each image, substantially reducing the total processing time.

### v3

@@ Abstract
[inst1]organization=Research Center of the Artificial Intelligence Institute, and Lab of Machine Learning and Knowledge Representation, Innopolis University, 420500,city=Innopolis, country=Russia, s.ahmadiasl@innopolis.ru
@@ 1 Introduction
Similar to SVD, which is used to analyze matrices, tensor decompositions are mathematical tools that generalize matrix decompositions (e.g., SVD, PCA) to higher-order tensors. They aim to represent a tensor as a combination of simpler, lower-dimensional components. The most common tensor decompositions include: Canonical polyadic decomposition [16], Tucker decomposition [34], tensor train decomposition [29], tensor ring decomposition [40] and tensor SVD (T-SVD) [22]. Tensor decompositions have numerous applications across various domains, such as signal processing to machine learning, tensors are indispensable tools in modern data science. The randomized block Krylov subspace method was proposed in [26] and was shown to be a very efficient technique for the low-rank approximation of matrices. Motivated by these results, it was extended to several tensor decompositions. For example, it was previously utilized in [30] and [36] for Tucker and TT decompositions, respectively. It is known that the randomized block Krylov subspace method can better capture the range of a matrix compared to the classical randomized algorithms, and this leads to obtaining better low-rank approximations. The integration of randomization and Krylov methods is a powerful paradigm shift. Randomization provides a cheap, parallelizable way to guess a good subspace, while Krylov methods provide a powerful, iterative way to refine that guess to high accuracy. The synergy between them creates algorithms that are faster, more robust, and come with stronger theoretical guarantees than either approach used in isolation. This has made them the method of choice for large-scale low-rank matrix approximations in both academic research and industrial practice. It is worth highlighting that randomized algorithms are applicable for computing low-rank approximations of data tensors that have low-rank structures, such as images and videos.
In this paper, we propose to use the randomized block Krylov method for the computation of the truncated Tensor SVD (T-SVD). It is another extension of the method proposed in [26] to tensors. To the best of our knowledge, this is the first paper discussing the possibility of using the randomized block Krylov method for the T-SVD model. A detailed theoretical analysis of the proposed algorithm is discussed, and they are supported by simulation results. An application of the proposed randomized method to image completion is given. Our simulation results show the effectiveness of the proposed randomized algorithm for the synthetic and real-world datasets.
This work makes the following contributions:
  • 1.
    We introduce a randomized block Krylov method to efficiently approximate the truncated T-SVD.
  • 2.
    We present a detailed theoretical analysis, establishing guarantees for the accuracy of the proposed algorithm.
  • 3.
    Through extensive simulations, we validate our theoretical findings and demonstrate the practical utility of the method on image completion problems.
@@ 2 Preliminary concepts
This section presents some definitions and lemmas that will be used in the sequel. We denote tensors with calligraphic letters and matrices with capital letters. The set of real and natural numbers denote by $\mathbb{R}$ and $\mathbb{N}$, respectively. Let $O$, and $\mathcal{O}$ stand for a zero matrix, and a zero tensor, respectively. The identity matrix of size $n\times n$ is denoted by $I_{n}$.
@@
where $\mathcal{U}\in\mathbb{R}^{n_{1}\times n_{1}\times n_{3}}$ is an orthogonal tensor, $\mathcal{S}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$ is an f-diagonal tensor (i.e., each frontal slice is a diagonal matrix), $\mathcal{V}\in\mathbb{R}^{n_{2}\times n_{2}\times n_{3}}$ is an orthogonal tensor. The number of nonzero tubes in the middle tensor $\mathcal{S}$ is called the tubal rank of $\mathcal{X}$ and is denoted by TrankT-rank. The nonzero tubes of the tensor $\mathcal{S}$, are called singular tubes. The truncated T-SVD of tubal rank $R$ for the tensor $\mathcal{X}$} is
@@ 1 $\overline{{\mathcal{X}}}=\mathtt{fft}\left({{\mathcal{X}},[],3}\right)$;
1 $\overline{{\mathcal{X}}}=\mathtt{fft}\left({\mathcal{X}},[],3\right)$;
@@ 3 Proposed randomized block Krylov method for approximate truncated T-SVD
The randomized block Krylov subspace algorithm is an efficient method for approximating large-scale matrix computations, including eigenvalue problems and low-rank matrix approximations. It combines randomization techniques with the power of Krylov subspace methods to achieve high accuracy with reduced computational cost. In the case of matrices as second-order tensors, it is known that the convergence rate of pure Krylov methods depends on the spectral gap. Slow decay (small gap) leads to slow convergence. Randomized sketching with power iterations creates a matrix $(XX^{T})^{q}X$ whose singular values decay faster, effectively enhancing the gap. Randomized methods provide strong, non-asymptotic probabilistic error bounds. For a target rank $k$, the approximation error is guaranteed (with high probability) to be within a small factor of the optimal error. Krylov refinement tightens these bounds.
Although a classical randomized algorithm applies power iterations $q\geq 1$ to enhance the decay of singular tubes, it does not use the results from each step of the iteration. Since it only uses the final result, $(\mathcal{X}*\mathcal{X}^{\top})^{q}*\mathcal{X}*\mathcal{B}$, the method generally cannot closely approximate the largest $k$ singular tubes. However, the block randomized algorithm captures a much richer spectral range, leading to faster, super-linear convergence and higher accuracy for the same number of tensor operations. Let us discuss this precisely. The classical randomized algorithm with the power iteration technique for truncated T-SVD is presented in Algorithm 4. As can be seen, it uses the last iteration $(\mathcal{X}*\mathcal{X}^{\top})^{q}*\mathcal{X}*\mathcal{B}$ of the power iteration loop for computing an orthonormal basis of the tensor range in Line 7. The proposed randomized algorithm looks for a better matrix to capture the range of the matrix and do this, it uses all iterations $(\mathcal{X}*\mathcal{X}^{\top})^{i}*\mathcal{X}*\mathcal{B},\,\,i=1,2,\ldots,q$, which leads to better capturing the range of the tensor and better approximation results. This process is summarized in Algorithm 5. The main difference between the proposed randomized algorithm and the one proposed in [37] lies in how they compute an orthonormal basis for the range of a tensor to be used in the projection stage within the randomization framework. We extensively examine this idea with detailed theoretical analyses.
It is worth noting that the primary trade-off is between accuracy and computational cost per iteration:
  • 1.
    Classical Randomized (with Power Iteration): Lower cost per iteration. However, it requires more iterations (higher q) to achieve high accuracy, especially for complex spectra, resulting in a high total cost for accurate approximations.
  • 2.
    Proposed Block Krylov: Higher cost per iteration due to the orthogonalization of a larger block Krylov basis. The significant advantage is that it achieves much higher accuracy for the same number of iterations “q”. Therefore, to achieve a high target accuracy, the Block Krylov method often has a lower total computational cost as it requires far fewer iterations (“q”). This trade-off is favorable for applications requiring high-precision approximations.
@@ 9 $\widehat{\mathcal{U}},{\mathcal{S}},{\mathcal{V}}]={\rm Truncated}$ ${\rm T-SVD}({\mathcal{B}},R)$;
Output : Truncated SVD: ${\mathcal{X}}\cong\widehat{\mathcal{U}}_{R}*\widehat{\mathcal{S}}_{R}*\widehat{\mathcal{V}}^{T}_{R}$
@@ 11 $\widehat{\mathcal{U}}=\mathcal{Q}*\mathcal{U}_{c}$;
$\widehat{\mathcal{U}}_{R}=\widehat{\mathcal{U}}(:,1:R,:),\,\widehat{\mathcal{S}}_{R}=\mathcal{S}_{c}(1:R,1:R,:),\,\widehat{\mathcal{V}}_{R}=\mathcal{V}_{c}(:,1:R,:).$
@@ 4 Theoretical results
In this section, we present a detailed theoretical analysis of the proposed randomized subspace block Krylov method. A key point that needs to be highlighted is that the proposed algorithm indeed applies the randomized block Krylov method on each frontal slice of the tensor in the Fourier domain. This facilitates the derivation of the upper bound approximation obtained by the proposed algorithm.
@@
Since the spectral norm on matrices is unitarily invariant and $\mathtt{bcirc}$ of an orthogonal tensor is an orthogonal matrix, we have
@@
Let $\mathcal{A}\in\mathbb{R}^{n\times n\times n_{3}}$ be a T-symmetric tensor and $\mathcal{C}$ be a real-valued tensor with adequate dimension. If $\mathcal{A}$ is a T-positive semidefinite tensor, then $\mathcal{C}^{\top}*\mathcal{A}*\mathcal{C}$ is a T-positive semidefinite tensor.
@@
Since each term of $\mathcal{D}$ is T-positive semidefinite, it can be easily shown that $\mathcal{D}$ is also semidefinite. So, from Lemma 11, we have
@@
where $\tilde{\mathcal{U}}$ and $\tilde{\mathcal{V}}$ are orthogonal tensors and $\tilde{\mathcal{S}}$ is an f-diagonal tensor, in which the entries on the first frontal slice are the singular values of the tensor $\mathcal{S},(\mathcal{S}*\mathcal{S}^{\top})*\mathcal{S},\ldots,(\mathcal{S}*\mathcal{S}^{\top})^{q}*\mathcal{S}]$. In other words, they are the square root of the eigenvalues of the following tensor
@@ 5 Experimental results
The proposed randomized algorithm and the baseline (Algorithm 4) are applied to the mentioned data tensor with the parameters P=5,q=2P=5,\,q=2 and $R=45$. The relative errors and running times of the algorithms are reported in Figures 1-3. We should point out that in Algorithm 5, in Line 7, the factor $\mathcal{Q}$ was truncated to $k+p$ term as our simulations showed that there is no difference between truncation and the full factor for this synthetic data, and our reported running times are for this case. If we consider the full factor $\mathcal{Q}$, the running time of Algorithm 5 will be higher than Algorithm 4. What we see from our numerical experiments is that the proposed randomized algorithm can provide estimates with better accuracy compared to the baseline, with a slightly higher running time. As was discussed in Section 3, the superior performance of Algorithm 5, is because it can capture a much richer spectral range, leading to faster convergence and higher accuracy for the same number of tensor operations.
We also made a sensitivity analysis on the oversampling and power iteration parameters. Our observation was that for tensors with frontal slices with fast singular value decay, the small oversampling parameter, e.g. $p=5,10,20$ is sufficient, often independent of the target rank $R$. Additionally, for the power iteration parameter, small integers, e.g., $q=0,1,2,3$, were sufficient and rarely exceeded $q=4$ due to numerical stability concerns.
@@
Figure 3: Relative error and computing time for Algorithms 4 and 5 for Example 1, Case 3.
@@
(Image compression) In this example, we consider the application of the proposed randomized block Krylov subspace method for compressing color images. We consider three color images “peppers”, “baboon” and “house” depicted in Figure 4. All images have size $256\times 256\times 3$. Then, the proposed randomized approach and the benchmark Algorithm 4 were applied to the three images using P=5,q=2P=5,\,q=2 and the tubal rank $R=25$. The reconstructed images computed by them are shown in Figure 4. The proposed algorithm can achieve marginally better results than the baseline. At the same time, its running time is a little higher due to dealing with a block of matrices (the tensor $\mathcal{K}$ in Algorithm 5) instead of only one single matrix (the tensor $\mathcal{B}$ in Line 7 of Algorithm 4). This requires extra computations. The simulation results indicate that the proposed randomized algorithm yields better results than the baseline. The sensitivity analysis on the oversampling and power iteration parameters was the same as the first example. More precisely, for the oversampling parameter, we achieved smooth improvement for 1 to 10. For the power iteration from $q=1$ to, $q=2,$ we saw a significant improvement, while for the $q\geq 2$ the improvement was not so much.
@@
(Image completion) In this experiment, we apply the proposed randomized block Krylov subspace method for the data completion task. We use the methodology proposed in [1, 4], where a low-rank approximation is required in the completion process. To ensure the paper is self-contained, we will briefly describe it. The image completion can be formulated as an optimization problem:
@@
${{P}_{\mathcal{T}}}\left({\mathcal{X}}\right)=\left\{\begin{array}[]{l}{x_{{i_{1}},{i_{2}},\ldots,{i_{N}}}}\,\,\,\,\,\,\left({{i_{1}},{i_{2}},\ldots,{i_{N}}}\right)\in\mathcal{T}.\\ 0\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm Otherwise}.\end{array}\right$.
@@
${{P}_{{\mathcal{T}}}}({{\mathcal{C}}})={{P}_{{\mathcal{T}}}}({{\mathcal{M}}})$ (27)
@@
Three completion methods as baselines, including: Bayesian tensor CPD (Canonical Polyadic decomposition) completion [^3] https://github.com/qbzhao/BCPF [39], tensor ring completion [^4] https://github.com/HuyanHuang/Tensor-completion-via-tensor-ring-decomposition [18] and tensor train completion [^5] https://github.com/zhaoxile/ [12] were used. The comparison was made for the “kodim03” image and for high missing ratios. The recovered images by these algorithms are shown in Figure 9. For all of these algorithms, the default parameters and tensor ranks are used in the demo files of their GitHub pages. Our experiments show the method is highly effective for image completion, performing robustly under both random and structured missing patterns. Comparing our results with the baselines for 95%, and 98% shows significant improvement. The Bayesian tensor CPD and tensor train completion were very slow, but the tensor ring completion algorithm was faster. The running time of our algorithm was comparable to the tensor ring completion, but it was approximately 4-5 times faster. For a fair comparison, we also included two randomized tensor completion methods [2, 3]. Using the same baseline images—Kodim01, Kodim02, and Kodim03—we reshaped them into tensors of size $8\times 8\times 8\times 8\times 8\times 4\times 9$ and applied a tensor ring rank of $(10,10,10,10,10,10,10)$. With the method from [2], we obtained PSNRs of $(27.45\,{\rm dB},\,22.65\,{\rm dB},\,28.11\,{\rm dB})$ [^6] The first, second, and third components are the results for Kodim01, Kodim02, and Kodim03, respectively. and computation times of (30.34, 31.49,(30.34,\,31.49,  29.98)\,29.98) [^7] The first, second, and third components are the results for Kodim01, Kodim02, and Kodim03, respectively.(in seconds). The method in [3] achieved PSNRs of $(29.94\,{\rm dB},$  $24.41\,{\rm dB},\,30.19\,{\rm dB})$ and computation times of (12.56,13.88, 13.22)(12.56\,,13.88,\,13.22)(in seconds). Our approach produced results very similar to those of [3] while being slightly faster. Moreover, compared to the method in [1], our algorithm was significantly faster and more efficient. These findings further confirm that the proposed randomized algorithm is highly effective for completion tasks.
@@
Figure 9: The completion results obtained by a) Bayesian tensor CPD (Canonical Polyadic decomposition) completion [39], b) tensor ring completion [18], and c) tensor train completion [12] for images with high missing ratios, for example 3.
@@
For this experiment, we use the Kodak dataset [^8] https://r0k.us/graphics/kodak/, comprising 23 images of size either $768\times 512\times 3$ or $512\times 768\times 3$. For consistency, all images were resized to $256\times 256\times 3$. We remove 80% of pixels in all images and use tubal tensor rank $R=40$, the oversampling parameter, $P=10$, and the power iteration parameter $q=2$. We use 100 iterations of the completion algorithm for the image recovery process. The PSNRs of the recovered images for the whole Kodak dataset are shown in Figure 10. This experiment demonstrated the feasibility of the completion algorithm, which is empowered by the proposed randomized block Krylov method. On average, completion of each image took 13 seconds; however, the algorithm’s computation can be significantly accelerated on GPU devices. Furthermore, when processing a batch of images, the completion process can be executed in parallel for each image, substantially reducing the total processing time.